How do I solve this: 3 over z = 9 over 5..............? (by "over" I mean 3 and 9 are numerators........z and 5 are denominators....
Solving for "z"...........Thank you

Answer by stanbon(48546)

solve this: 3 over z = 9 over 5
3/z = 9/5
Invert both sides to get:
z/3 = 5/9
Multiply both sides by 3 to get;
z = 5/3
===========
Cheers,
Stan H.